George is five times as old as Jane.
1) Let x stand for Jane's age and y stand for George's age. Write an equation in terms of x and y to represent the above statement.
2) In fifteen years time George will be three times as old as Jane. Use this relation to show that y = 3x + 30.
3) Use the equations obtained in the previous parts of the question to determine the values of x and y.

Answer by mananth(16946)

y=5x
after 15 years
y+15=3(x+15)
y+15=3x+45
y= 3x+45-15
y=3x+30
substitute y=5x
5x=3x+30
2x=30
/2
x=15
y=5x
y=5*15
y=75